Crasper & Bros, Inc. manufactures phone cases for both Androids and iPhones. Last year, Crasper manufactured 34,060 units and sold 28,600 phone cases (units). They have reported the following Production costs for the year:

Direct materials $ 255,450
Direct labor $ 173,706
Variable manufacturing overhead $ 269,074
Fixed manufacturing overhead $ 510,900

Sales of their phone cases totaled $1,172,600 for the year, variable selling and administrative expenses totaled $148,720, and fixed selling and administrative expenses totaled $214,578. Since they are a brand new company, there was no beginning inventory. Assume that the direct labor cost is a variable cost.

Using the absorption costing, the ending inventory for the year would be valued at:

Multiple Choice

  • $263,830

  • $255,330

  • $193,830

  • $221,330
Answer #1

Answer: $193,830

Unit Product Cost Under Absorption Costing

Direct Material 255,450
Direct Labor 173,706
Variable MOH 269,074
Fixed MOH 510,900
Total Product Cost for 34,060 Units 1,209,103
Units Product Cost =1209,130/34,060 $35.50

Ending inventory in Units = Prodcued -Sold =34,060-28,600 = 5,460

Ending inventory in dollars =Ending Units * Product Cost = 5,460*35.50= $193,830(Answer)